OpenStudy (danielbarriosr1):

Ms. Jones wants to create a blend of coffee. one coffee is priced at $8.40 per pound and the other is $7.28 per pound. How many pounds of the $7.28 per pound coffee should she mix with 9 pounds of the $8.40 per pound coffee to sell the mixture for $7.95 per pound? round the answer to the nearest hundredth of a pound.

OpenStudy (danielbarriosr1):

The possible answers are: A. 6.04 pounds B. 4.05 pounds C. 7.13 pounds D. 7.00 pounds

OpenStudy (danielbarriosr1):

Thanks anyway @KirbyLegs

OpenStudy (danielbarriosr1):

nop

OpenStudy (danielbarriosr1):

I dont get it

OpenStudy (danielbarriosr1):

can you explain me please?

OpenStudy (anonymous):

8.40*9+7.28x=7.95(9+x) 75.6+7.28x=71.55+7.95x 7.28x-7.95x=71.55-75.60 -.67x=-4.05 x=-4.05/-.67 x=6.045 pounds of $7.28 coffee is used. Proof: 8.40*9+7.28*6.045=7.95(9+6.045) 75.6+44=7.95*15.045 119.6=119.6

OpenStudy (anonymous):

snap, I reversed the variables, look at it and change the variables and solve using this example

OpenStudy (anonymous):

it should have been 8.40x and 7.28*9

OpenStudy (danielbarriosr1):

ok thanks now i get it